# Over-the-Air DPD and Reciprocity Calibration in Massive MIMO and Beyond

### v2

@@ Abstract
Non-linear transceivers and non-reciprocity of downlink and uplink channels are two major challenges in the deployment of massive multiple-input-multiple-output (MIMO) systems. We consider an over-the-air (OTA) approach for digital pre-distortion (DPD) and reciprocity calibration to jointly address these issues. In particular, we consider a memory-less non-linearity model for the base station (BS) transmitters, and we propose a method to linearize the transmitters and perform the calibration by using mutual coupling OTA measurements between BS antennas. We show that, by using only the OTA-based data, we can linearize the transmitters and design the calibration to compensate for both the non-linearity and non-reciprocity of BS transceivers effectively. This allows to alleviate the requirement to have dedicated hardware modules for transceiver characterization. Moreover, the proposed reciprocity calibration method is solely based on closed-form linear transformations, achieving a significant complexity reduction over state-of-the-art methods, which usually rely on costly iterative computations. Simulation results showcase the potential of our approach in terms of the calibration matrix estimation error and downlink data-rates when applying zero-forcing (ZF) precoding after using our OTA-based DPD and reciprocity calibration method.
@@ I Introduction
Massive multiple-input multiple-output (MIMO) has been one of the main technologies in the development of the fifth-generation (5G) of wireless networks, by enabling significant improvements in network capacity and reliability [1, 2]. In the early stages of massive MIMO development, several proposals motivated the adoption of frequency-division duplexing (FDD) in massive MIMO deployments. While some advantages may arise from considering FDD [3], the over-head in downlink channel estimation is an important drawback which limits the system scalability. Therefore, time-division duplexing (TDD) is selected as the more viable approach for the deployment of massive MIMO in 5G and beyond, since it enables downlink channel estimation based on uplink channel state information (CSI) and channel reciprocity [4].
@@
In this paper, we propose a method exploiting OTA measurements of inter-antenna mutual couplings at the BS to perform both the DPD and reciprocity calibration. The reciprocity methods in the literature mostly rely on linear transceivers, i.e., modeled by a complex number [6, 5]. This assumption is not accurate in practical cases, especially when scaling up massive MIMO systems, which necessitates deploying less expensive non-linear components and non-ideal linearization techniques for cost-efficiency reasons. Therefore, we assume that the TX-chains in the BS are non-linear and we propose a method to linearize them based on mutual coupling measurements. Our approach relies on the same hardware available for OTA-based reciprocity calibration, which improves resource efficiency by eliminating the need for dedicated hardware to perform per-antenna DPD. Considering the linearized transmitters after applying our OTA-based method, we derive a reciprocity calibration approach which depends only on linear closed-form transformations. Thus, the proposed reciprocity calibration allows for a significant complexity reduction over iterative methods as the ones considered in [6, 5]. Numerical results show that, although we consider non-linear TX-chains, the calibration matrix estimation error is fairly close to the Cramer-Rao Lower Bound (CRLB) derived in [6]. We also evaluate the system performance in downlink data transmission using zero-forzing (ZF) precoding and show that the proposed calibration method can approach the optimal calibration performance for moderate values of signal to noise ratio (SNR).
@@ II System Model
We consider a TDD multi-user-MIMO scenario where an $M$-antenna BS serves $K\leq M$ user equipments through a narrow-band channel. We assume that the UE transceivers operate in the linear regime during UL and DL. The BS receivers are also assumed to operate in the linear regime for the UL, but the BS transmitters exhibit non-linear response for the DL. [^1] The assumption of having non-linear behavior only in the BS transmitters is reasonable taking into account that this is where the input power is significantly higher, pushing the power amplifiers to the non-linear regime [11].
@@ II-A Uplink
where $\boldsymbol{R}_{\mathrm{B}}=\mathrm{diag}(r^{\mathrm{B}}_{1},\dots,r^{\mathrm{B}}_{M})$ and $\boldsymbol{T}_{\mathrm{U}}=\mathrm{diag}(t_{1}^{\mathrm{U}},\dots,t^{\mathrm{U}}_{K})$ are associated with the linear response of the BS receivers and the UE transmitters, respectively, and $\boldsymbol{H}$ corresponds to the $M\times K$ reciprocal propagation channel matrix [6]. Note that, if the UEs transmitters had non-linear behavior, the term $\boldsymbol{T}_{\mathrm{U}}$ in the estimated UL channel would be substituted by a non-linear function of the pilot matrix. A thorough study of its implications may be studied in future work, but the presented method would still be able to cope with the non-linearity and non-reciprocity originated at the BS side.
The UL channel, which includes the effects of the UE transmitters and the BS receivers, is assumed to be perfectly estimated at the BS based on UL pilots transmitted by the UEs. Hence, perfect knowledge of $\boldsymbol{H}_{\mathrm{UL}}$ is available at the BS, allowing for effective implementation of linear processing techniques, e.g., ZF, linear minimum mean squared error (LMMSE), and maximum ratio combining (MRC), to estimate $\boldsymbol{s}_{\mathrm{U}}$. However, the diagonal entries of $\boldsymbol{T}_{\mathrm{U}}$ and $\boldsymbol{R}_{\mathrm{B}}$, which may take arbitrary complex numbers, are unknown since they can be affected by parameters such as temperature, hardware imperfections, etc.
@@ II-B Downlink
where $\boldsymbol{x}_{\mathrm{B}}$ is the $M\times 1$ vector of input symbols to each BS TX-chain, $\boldsymbol{n}_{\mathrm{U}}\sim\mathcal{CN}(\boldsymbol{0},\mathrm{diag}(N_{0,\mathrm{U}_{1}},\dots,N_{0,\mathrm{U}_{K}}))$ models the AWGN at the UEs, $\boldsymbol{R}_{\mathrm{U}}=\mathrm{diag}((r^{\mathrm{U}}_{1},\dots,r^{\mathrm{U}}_{K})$ is associated with the linear response of the UE receivers, while $\boldsymbol{f}:\mathbb{C}^{M\times 1}\rightarrow\mathbb{C}^{M\times 1}$ is a vector-valued function modeling the non-linear response of the BS TX-chains. We assume that the transmitted symbols are generated such that
@@
where $\boldsymbol{s}_{\mathrm{B}}$ is the $K\times 1$ vector of symbols intended for the UEs, $\boldsymbol{W}$ is the $M\times K$ linear precoding matrix applied at the BS baseband unit (BBU), while $\boldsymbol{g}:\mathbb{C}^{M\times 1}\rightarrow\mathbb{C}^{M\times 1}$ is the non-linear vector-valued function associated to the DPD applied at each TX-chain.
Let us assume that the cross-talk between TX-chains is negligible so that $\boldsymbol{f}(\cdot)$, and correspondingly $\boldsymbol{g}(\cdot)$, are component-wise functions. Considering a third-order memory-less polynomial model [12], we express
@@
where $t^{\mathrm{B}}_{m}$ and $\beta_{m}$ are two complex scalars characterizing the non-linear response of the $m$’th TX-chain at the BS. The reason for focusing on the third-order non-linearity model is that it allows capturing the main idea of our solution, while maintaining a neat exposition. However, the main results of this work have trivial extension for higher order models, as remarked in the following section. Alternatively, one can fit any RF non-linear behavior to the 3rd order model, which should still capture its main impact [13]. In general, the BS TX-chain and UEs RX-chain responses are unknown, which means that the non-linear parameters $t^{\mathrm{B}}_{m}$ and $\beta_{m}$, as well as the diagonal entries of $\boldsymbol{R}_{\mathrm{U}}$, are unknown at the BS. Note that, unlike traditional work on reciprocity calibration [6], where $\boldsymbol{f}(\cdot)$ is associated to a linear transformation $\boldsymbol{T}_{B}=\mathrm{diag}(t^{\mathrm{B}}_{1},\dots,t^{\mathrm{B}}_{M})$, we cannot hereby define an aggregated DL channel matrix due to the non-linear nature of the TX-chains.
@@ II-C Background: OTA Reciprocity Calibration
As mentioned earlier, previous work has addressed the problem of reciprocity calibration in massive MIMO assuming BS TX-chains that operate in the linear regime [6, 5]. Under such assumptions, we may define the DL channel matrix as
@@
which may be also derived from the presented system model, assuming $\beta_{m}=0$ in (5). [^2] Equivalently, the reciprocity calibration problem from [6] may be obtained by assuming perfect DPD up to unknown scalars, i.e., $\boldsymbol{f}(\boldsymbol{g}(\boldsymbol{x}))=\boldsymbol{T}_{\mathrm{B}}\boldsymbol{x}$. The main goal of reciprocity calibration methods, as the one in [6], is to estimate the reciprocity matrix, defined as
@@
The reason is that, if we have knowledge of $\boldsymbol{C}$, we can transform the known UL channel matrix into
@@
Note that $\widetilde{\boldsymbol{H}}_{\mathrm{DL}}$ corresponds to $\boldsymbol{H}_{\mathrm{DL}}$ up to an unknown $K\times K$ diagonal matrix, namely $\boldsymbol{D}=\boldsymbol{T}_{\mathrm{U}}\boldsymbol{R}_{\mathrm{U}}^{-1}$, multiplied from the left. Hence, $\widetilde{\boldsymbol{H}}_{\mathrm{DL}}$ can be effectively used for linear precoding, with the only caveat that the symbols received by the UEs would end up multiplied by an unknown scalar, which has negligible impact on system performance [14]. [^3] In the practical testbed considered in [6], this issue is addressed by sending an extra DL pilot. We may thus ignore the non-reciprocity associated to the UEs hardware, modeled by $\boldsymbol{R}_{\mathrm{U}}$ and $\boldsymbol{T}_{\mathrm{U}}$, and focus on characterizing the non-reciprocity associated to the BS.
An important advantage of the OTA-based calibration methods which are based on mutual coupling measurements is that they avoid the need for dedicated hardware to characterize the linear response of each transceiver chain, and can improve the cost-efficiency of massive MIMO systems [6, 5]. Similarly, we can argue that having dedicated hardware to perform DPD may compromise the cost-efficiency of MIMO systems with increasing number of antennas, e.g., massive MIMO and beyond. Thus, we next propose a method to jointly characterize the non-linear response of the BS TX-chains, as well as the resulting reciprocity matrix, to suitably design $\boldsymbol{g}(\cdot)$ and $\boldsymbol{W}$ for effectively serving the UEs in the DL.
@@ III OTA DPD and reciprocity calibration
Our proposed OTA-based solution may be divided into three stages:
@@
  • Third, reciprocity calibration is performed based on the DPD-linearized BS TX-chains, after which effective DL precoding, associated to $\boldsymbol{W}$, would become available at the BS.
@@ III-A OTA non-linearity characterization
where $h_{ij}$ is the mutual coupling gain between antennas $i$ and $j$, which is assumed fixed and known at the BS, [^4] The coupling gains may be characterized with a single complete measurement of the antenna system using a network analyzer [6]. Thus, knowledge of these may be assumed in any MIMO-related scenario with co-located TX antennas. Nevertheless, our methods can be easily extended to the case where these coefficients are unknown, as outlined throughout this work. $x_{i,\ell}$ is the $\ell$’th pilot symbol transmitted by the $i$’th antenna, and $n_{ij,\ell}\sim\mathcal{CN}(0,N_{0})$ models the measurement noise. Note that we have removed the superscript B from the parameters $r_{j}$ and $t_{i}$ for notation convenience since, as previously reasoned, we may focus on the non-reciprocity associated to the BS.
For each pair of non-linearity parameters associated to one TX-chain, there are $M-1$ relevant DPD measurements per pilot transmission, i.e., all of those originated in the same antenna, but received at different antennas. Thus, each of these measurements would share the same $t_{i}$ and $\beta_{i}$ in (9), but they would be related to a different complex gain $r_{j}$, associated to the linear response of RX-chain from the respective receiving antenna. Since the complex gains $r_{j}$ are unknown, it is not possible to directly estimate the non-linearity parameters $t_{i}$ and $\beta_{i}$ from this dataset. However, we may combine the $M-1$ measurements by averaging them after compensating for the known mutual coupling gains, so as to reduce the uncertainty, as well as the resulting noise. The combined measurements are then given by
@@
Note that, in case of considering non-linearity polynomial models of higher order, the vector $\bm{\theta}$ (correspondingly $\bm{\Phi}$) would include one term per polynomial coefficient [10]. However, the presented method would still be applicable.
@@ III-B DPD linearization
In this stage the non-linearity parameters estimated in the previous stage are used to linearize the output via DPD, i.e., by adjusting $\boldsymbol{g}(\cdot)$ in (4). Let us assume that $\bm{\theta}_{i}$ has been perfectly estimated from (12). The true non-linearity to compensate is the 3rd-order nonlinear function given in (5). However, the estimated non-linearity parameters in (14) characterize a different component-wise function given by
@@
Since the function $\tilde{\boldsymbol{f}}(\cdot)$ is fully characterized, we can find its inverse by using methods such as the postdistortion approach [16]. We may then select
@@
which is applied to the the transmitted symbols as described in (4). The resulting symbols transmitted through the reciprocal channel, which are given by $\boldsymbol{f}(\boldsymbol{x}_{\mathrm{B}})$ in (3), may then be expressed as
@@
Hence, applying the proposed OTA-DPD, results in an equivalent linear transmitter gain given by $\tilde{\boldsymbol{T}}_{\mathrm{B}}=\boldsymbol{Q}^{-1}$. Now that the transmitter is linear with an unknown complex gain, we can define a DL channel matrix equivalent to (6), but substituting $\boldsymbol{T}_{\mathrm{B}}$ for $\tilde{\boldsymbol{T}}_{\mathrm{B}}$, so that reciprocity calibration methods as those presented in [6, 5] are directly applicable. However, we will show that the reciprocity calibration can be performed without the need for complex iterative methods.

### v3

@@ Abstract
Non-linear transceivers and non-reciprocity of downlink and uplink channels are two major challenges in the deployment of massive multiple-input-multiple-output (MIMO) systems. We consider an over-the-air (OTA) approach for digital pre-distortion (DPD) and reciprocity calibration to jointly address these issues. In particular, we consider a memory-less non-linearity model for the base station (BS) transmitters, and we propose a method to perform both linearization and reciprocity calibration based on mutual coupling OTA measurements between BS antennas. We show that, by using only the OTA-based data, we can linearize the transmitters and design the calibration to compensate for both the non-linearity and non-reciprocity of BS transceivers. This allows alleviating the requirement to have dedicated hardware modules for transceiver linearization. Moreover, the proposed reciprocity calibration method is solely based on closed-form linear transformations, achieving a significant complexity reduction over state-of-the-art reciprocity methods, which assume linear transceivers, and rely on iterative methods. Simulation results showcase the potential of our approach in terms of the calibration matrix estimation error and downlink data-rates when applying zero-forcing (ZF) precoding after using our OTA-based DPD and reciprocity calibration method.
@@ I Introduction
Massive multiple-input multiple-output (MIMO) has been one of the main technologies in the development of the fifth-generation (5G) of wireless networks, by enabling significant improvements in network capacity and reliability [1]. In the early stages of massive MIMO development, several proposals motivated the adoption of frequency-division duplexing (FDD) in massive MIMO deployments. While some advantages may arise from considering FDD [2], the over-head in downlink channel estimation is an important drawback which limits the system scalability. Therefore, time-division duplexing (TDD) is selected as the more viable approach for the deployment of massive MIMO in 5G and beyond, since it enables downlink channel estimation based on uplink channel state information (CSI) and channel reciprocity [3].
@@
In this paper, we propose a method exploiting OTA measurements of inter-antenna mutual couplings at the BS to perform both the DPD and reciprocity calibration. The literature on reciprocity calibration has focused on the latter by assuming linearized transceivers [5, 4, 6, 7]. This assumption is not accurate in practical cases, especially when scaling up massive MIMO systems, which necessitates deploying less expensive non-linear components and non-ideal linearization techniques for cost-efficiency reasons. Therefore, we assume that the TX-chains in the BS are non-linear and we propose a method to linearize them based on mutual coupling measurements. Our approach relies on the same hardware available for OTA-based reciprocity calibration, which improves resource efficiency by eliminating the need for dedicated hardware to perform per-antenna DPD. Considering the linearized transmitters after applying our OTA-based method, we derive a reciprocity calibration approach which depends only on linear closed-form transformations. Thus, the proposed reciprocity calibration allows for a significant complexity reduction over iterative methods focusing on the reciprocity calibration problem [5, 4, 6, 7]. Numerical results show that, even with the extra challenge of having to deal with non-linear TX-chain compensation, the calibration matrix estimation error approaches the Cramer-Rao Lower Bound (CRLB), which was derived in [5]. We also evaluate the system performance in downlink data transmission using zero-forzing (ZF) precoding and show that the proposed method can approach the perfect calibration performance even though we are considering non-linear TX-chains.
@@ II System Model
We consider a TDD multi-user-MIMO scenario where an $M$-antenna BS serves $K\leq M$ user equipments through a narrow-band channel. For the UEs, we assume that the TX- and RX-chains are both operating in the linear regime. For the BS, we assume that the RX-chains are also operated in the linear regime, but the TX-chains exhibit non-linear response. [^1] Assuming non-linear behavior only in the BS TX-chains is reasonable taking into account that this is where the input power is significantly higher, pushing the power amplifiers to the non-linear regime [10]. For the UEs, this non-linearity may be compensated with a single DPD module per UE.
@@ II-A Uplink
where $\boldsymbol{R}_{\mathrm{B}}=\mathrm{diag}(r^{\mathrm{B}}_{1},\dots,r^{\mathrm{B}}_{M})$ and $\boldsymbol{T}_{\mathrm{U}}=\mathrm{diag}(t_{1}^{\mathrm{U}},\dots,t^{\mathrm{U}}_{K})$ are associated with the linear response of the BS receivers and the UE transmitters, respectively, and $\boldsymbol{H}$ corresponds to the $M\times K$ reciprocal propagation channel matrix [5]. The UL channel in (2), which includes the effects of the UE transmitters and the BS receivers, can be estimated at the BS based on UL pilots transmitted by the UEs, allowing for effective implementation of linear processing techniques, e.g., ZF and maximum ratio combining (MRC).
Note that, if the UEs transmitters had non-linear behavior, the term $\boldsymbol{T}_{\mathrm{U}}$ in the estimated UL channel would be substituted by a non-linear function of the pilot matrix. A thorough study of this case may be considered in future work, but the presented method would still be able to cope with the non-linearity and non-reciprocity originated at the BS side.
@@ II-B Downlink
where $\boldsymbol{x}_{\mathrm{B}}$ is the $M\times 1$ vector of input symbols to each BS TX-chain, $\boldsymbol{n}_{\mathrm{U}}\sim\mathcal{CN}(\boldsymbol{0},\mathrm{diag}(N_{0,\mathrm{U}_{1}},\dots,N_{0,\mathrm{U}_{K}}))$ models the AWGN at the UEs, $\boldsymbol{R}_{\mathrm{U}}=\mathrm{diag}((r^{\mathrm{U}}_{1},\dots,r^{\mathrm{U}}_{K})$ is associated with the linear response of the UE receivers, and $\boldsymbol{f}:\mathbb{C}^{M\times 1}\rightarrow\mathbb{C}^{M\times 1}$ is a vector-valued function modeling the non-linear response of the BS TX-chains. We assume that the transmitted symbols are generated such that
@@
where $\boldsymbol{s}_{\mathrm{B}}$ is the $K\times 1$ vector of symbols intended for the UEs, $\boldsymbol{W}$ is the $M\times K$ linear precoding matrix applied at the BS baseband unit (BBU), and $\boldsymbol{g}:\mathbb{C}^{M\times 1}\rightarrow\mathbb{C}^{M\times 1}$ is the non-linear vector-valued function associated to the DPD applied at each TX-chain.
Let us assume that the cross-talk between TX-chains is negligible so that $\boldsymbol{f}(\cdot)$, and correspondingly $\boldsymbol{g}(\cdot)$, are component-wise functions. Considering a third-order memory-less polynomial model [11], we have
@@
where $t^{\mathrm{B}}_{m}$ and $\beta_{m}$ are two complex scalars characterizing the non-linear response of the $m$’th TX-chain at the BS. In general, the BS TX-chain and UEs RX-chain responses are unknown, which means that the non-linear parameters $t^{\mathrm{B}}_{m}$ and $\beta_{m}$, as well as the diagonal entries of $\boldsymbol{R}_{\mathrm{U}}$, are unknown at the BS. Note that, unlike state-of-the-art work on reciprocity calibration [5, 4, 6, 7], where $\boldsymbol{f}(\cdot)$ is associated to a linear transformation $\boldsymbol{T}_{B}=\mathrm{diag}(t^{\mathrm{B}}_{1},\dots,t^{\mathrm{B}}_{M})$, we cannot hereby define an aggregated DL channel matrix due to the non-linear nature of the TX-chains.
@@ II-C Background: OTA Reciprocity Calibration
As mentioned earlier, previous work has addressed the problem of reciprocity calibration in massive MIMO assuming BS TX-chains operating in linear regime [5, 4, 6, 7]. Under such assumptions, we may define the DL channel matrix as
@@
which may be also derived from the presented system model, assuming $\beta_{m}=0$ in (5). [^2] Equivalently, reciprocity calibration problems in [5, 4, 6, 7] may be obtained by assuming perfect DPD up to unknown scalars, i.e., $\boldsymbol{f}(\boldsymbol{g}(\boldsymbol{x}))=\boldsymbol{T}_{\mathrm{B}}\boldsymbol{x}$. The main goal of reciprocity calibration methods is to estimate the reciprocity matrix,
@@
The reason is that, if we have knowledge of $\boldsymbol{C}$, we can transform the estimated UL channel matrix into
@@
Note that $\widetilde{\boldsymbol{H}}_{\mathrm{DL}}$ corresponds to $\boldsymbol{H}_{\mathrm{DL}}$ up to an unknown $K\times K$ diagonal matrix, namely $\boldsymbol{D}=\boldsymbol{T}_{\mathrm{U}}\boldsymbol{R}_{\mathrm{U}}^{-1}$, multiplied from the left. Hence, $\widetilde{\boldsymbol{H}}_{\mathrm{DL}}$ can be effectively used for linear precoding, with the only caveat that the symbols received by the UEs would end up multiplied by an unknown scalar, which has negligible impact on system performance [12]. [^3] In practice, this issue is addressed by sending a DL pilot [5]. We may thus ignore the non-reciprocity associated to the UEs hardware, modeled by $\boldsymbol{R}_{\mathrm{U}}$ and $\boldsymbol{T}_{\mathrm{U}}$, and focus on characterizing the non-reciprocity associated to the BS. An important advantage of the OTA-based calibration methods which are based on mutual coupling measurements is that they avoid the need for dedicated hardware to characterize the linear response of each TX-chain, and can improve the cost-efficiency of massive MIMO systems [5]. Similarly, we can argue that having dedicated hardware to perform DPD may compromise the cost-efficiency of MIMO systems with increasing number of antennas, e.g., massive MIMO and beyond. Thus, we next propose a method to jointly characterize the non-linear response of the BS TX-chains, as well as the resulting reciprocity matrix, to suitably design $\boldsymbol{g}(\cdot)$ and $\boldsymbol{W}$ for effectively serving the UEs in the DL.
@@ III OTA DPD and reciprocity calibration
Our proposed method may be divided into three stages:
@@
  • Third, reciprocity calibration is performed based on the DPD-linearized BS TX-chains, after which effective DL precoding, associated to $\boldsymbol{W}$, becomes available at the BS.
@@ III-A OTA non-linearity characterization
where $h_{ij}$ is the mutual coupling gain between antennas $i$ and $j$, which is assumed fixed and known at the BS, [^4] The coupling gains may be characterized with a single measurement of the antenna system using a network analyzer [5]. Thus, knowledge of these may be assumed in any MIMO-related scenario with co-located TX antennas. $x_{i,\ell}$ is the $\ell$’th pilot symbol transmitted by the $i$’th antenna, and $n_{ij,\ell}\sim\mathcal{CN}(0,N_{0})$ models the measurement noise. Note that we have removed the superscript B from the parameters $r_{j}$ and $t_{i}$ for notation convenience since, as previously reasoned, we may focus on the non-reciprocity associated to the BS.
For each pair of non-linearity parameters associated to one TX-chain, there are M1M-1 relevant DPD measurements per pilot transmission, i.e., all of those originated in the same antenna, but received at different antennas. Thus, each of these measurements would share the same $t_{i}$ and $\beta_{i}$ in (9), but they would be related to a different complex gain $r_{j}$, associated to the linear response of the RX-chain from the respective receiving antenna. Since the complex gains $r_{j}$ are unknown, it is not possible to directly estimate the non-linearity parameters $t_{i}$ and $\beta_{i}$ from this dataset. However, we may combine the M1M-1 measurements by averaging them after compensating for the known mutual coupling gains, so as to reduce the uncertainty, as well as the resulting noise. The combined measurements are then given by
@@
Note that, while we have presented our method for a 3rd order non-linearity model (5), which is the main source of inter-modulation terms falling within the operating frequencies, the method can be generalized for higher order non-linearity models as well. In case of considering non-linearity polynomial models of higher order, the vector $\bm{\theta}$ (correspondingly $\bm{\Phi}$) would include one term per polynomial coefficient and the presented method would still be applicable. Alternatively, one can fit any RF non-linear behavior to the 3rd order model, which should still capture its main impact [14].
@@ III-B DPD linearization
In this stage the non-linearity parameters estimated in the previous stage are used to linearize the output via DPD, i.e., by adjusting $\boldsymbol{g}(\cdot)$ in (4). The true non-linearity to compensate is the nonlinear function given in (5). However, the estimated non-linearity parameters in (14), $\theta_{1i}$ and $\theta_{2i}$, characterize a different component-wise function given by
@@
[…]
@@
which is applied to the the transmitted symbols as described in (4). In practice, perfect DPD inversion may not be fully achievable, mainly due to limited DPD size and imperfect estimation of non-linearity parameters. We have considered imperfect inversion in the numerical results from Section IV.
The resulting symbols transmitted through the reciprocal channel, may then be expressed as
@@
Hence, applying the proposed OTA-DPD, results in an equivalent linear transmitter gain given by $\tilde{\boldsymbol{T}}_{\mathrm{B}}=\boldsymbol{Q}^{-1}$. Now that the transmitter is linear, we can define a DL channel matrix equivalent to (6), but substituting $\boldsymbol{T}_{\mathrm{B}}$ for $\tilde{\boldsymbol{T}}_{\mathrm{B}}$, so that reciprocity calibration methods as those presented in [5, 4] are directly applicable. However, we will show that the reciprocity calibration can be performed without the need for complex iterative methods.